# Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides

### v1

@@ III.4 Physical model for resistive switching based on point defects
As discussed in Section S1 of the SI, a correlation between the measured switching energy [14] and the computed adsorption energies for the common TMDs suggests that the adsorption energy plays an important role in the energetics of resistive switching based on transition metal dichalcognides. Previous studies, such as the one by Frey et al [43], have used a similar idea of using binding energies and formation energies to identify optimal defects for applications in neuromorphic and quantum information processing. Frey et al. [43] used random forest regression to predict adsorption energies and identify potential descriptors but did not fully delve into the defect chemistry that dictated trends in adsorption energy. This study emphasizes understanding the adsorption mechanism between 2D TMDs and various transition metal adsorbates, to further elucidate how the valency, periodicity, and electronic configuration of the metal affect the adsorption energetics.
IV Discussion
@@
We have shown that the adsorption energy is relatively insensitive to the chemical species of the substrate when considering commonly studied TMDs (MoS2, MoSe2, WS2, WSe2). Saliently, the trends in adsorption species may be explained using the d-band model, a model based on electronic structure descriptors most well-known in its application to catalysts. Hence, the switching energy is materials-dependent and that dependence can be rationalized. An earlier study on the adsorption of Li in transition metal dichachalcogenides for battery applications demonstrated that the work function can serve as a simple descriptor for explaining trends in Li adsorption[79]. In this study, we generalize this concept to adsorption energies of metal adatoms in low-dimensional materials. Thus, insights from electrochemistry and catalysis may be effectively cross-applied in resistive switching contexts, and can be applied more broadly to classes of substrates and adsorbates than previously demonstrated.
@@
Within and beyond resistive switching, defects in low-dimensional materials play a prominent role in next generation computing devices. Indeed, two-dimensional (2D) materials are ideal platforms for hosting quantum defects [80, 81]. An important implication of our study in rationalizing adsorption energies of point defect complexes in TMDs is the identification of electronic structure and atomic descriptors for both the substrate and electrode, which extends our ability to engineering designer defects for targeted applications. For practical device applications, we expect that the adsorption energy of transition metal adatoms plays a role in the switching energy (see also Figure S2). That is, a large (negative) adsorption energy corresponds to a greater stability of the adsorbate, which may be important for long-term operation. On the other hand, a low (near zero) adsorption energy corresponds to a low switching energy, which may contribute to more energy-efficient resistive switching.
@@ V Conclusion
We investigated the adsorption energies of various transition metal adsorbates on different monolayer TMD substrates, utilizing both computational methods and theoretical models to understand the underlying trends. We evaluated the adsorption energy of a metal adatom on a pre-existing anion vacancy for MoS2, MoSe2, WS2, and WSe2 and suggest these adsorption energies as proxy for investigating resistive switching energy. The adsorption energies displayed significant variation depending on the periodic group of the transition metal, with early transition metals (Hf, Zr, Ti) exhibiting the strongest adsorption energies, and late transition metals (Cu, Au, Ag, Zn, Cd, Hg) showing the weakest, while the middle transition metals exhibited moderate adsorption energies in between the early and late transition metals. We demonstrated these trends in the adsorption energy could be largely captured by trends in the d-band center with respect to the Fermi level, and as a descriptor can capture most of the trend in adsorption energies. COHP analysis provided further insights into the nature of bonding and anti-bonding interactions between the adsorbates and their nearest neighbor sulfur atoms in the TMDs, emphasizing the importance of the local environment surrounding the adsorption site in determining the adsorption energy. We used this understanding to perform compressed sensing on atom-based descriptors, which consequently does not require a DFT calculation. This SISSO analysis demonstrated promise in predicting absolute adsorption energies, using compound descriptors based off of the ionization energy, electronegativity, and atomic radii of the adsorbate and adsorption site.
Our findings underscore the importance of considering both the electronic structure and the local bonding environment in understanding and predicting adsorption behavior on TMD substrates. In conclusion, this study provides a framework for predicting defect energetics in TMD systems and helps to pave the way for rational materials selection in resistive switching. By integrating electronic structure analysis with techniques inspired by machine learning, we can gain deeper insights into the fundamental interactions governing adsorption and tailor materials for various technological applications in resistive switching, catalysis, sensing, and beyond.
VI Acknowledgements
@@ VII Author contributions
VIII Data Availability
The data supporting the findings of this study are openly available on Zenodo at DOI: 10.5281/zenodo.13892108, Ref [83]
Supplementary Information for
Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides
Brian H. Lee, Jameela Fatheema, Deji Akinwande‡, †, Wennie Wang¶, †
Texas Materials Institute, University of Texas at Austin, USA
Chandra Family Department of Electrical and Computer Engineering, University of Texas at Austin, USA
McKetta Department of Chemical Engineering, University of Texas at Austin, USA
Email: wwwennie@che.utexas.edu
S1 Connection between switching energy and adsorption energy
Figure S1: The dissociation-diffusion-adsorption (DDA) model describing the switching mechanism in a vertical MoS2 and gold memristor. The model outlines the transition between high resistance state (HRS) and low resistance state (LRS) as follows: (a) the initial state of the device, (b) dissociation of a gold adatom from the electrode, (c) weak adsorption of the gold adatom onto a pristine MoS2 site, and (d) strong adsorption of the gold adatom onto a sulfur-vacancy site.. Adapted from [14].
The switching energy in memristors can be defined as the energy required to ”switch” the material between a low-resistance state (LRS) and high-resistance state (HRS). Thus, it is expected that this switching energy is related to adsorption energies according to the dissociation-diffusion-adsorption (DDA) model (Figure S1). To demonstrate this correlation, we analyzed the I-V curves of various monolayer transition metal dichalcogenides (TMDs) exhibiting non-volatile resistive switching (NVRS), as measured by Ge et al.[14]. The switching power was estimated by numerically integrating the hysteretic I-V curves, specifically calculating the SET switching power by integrating over the SET process of the hysteresis. We report the correlation between the integrals of the I-V curves during the SET transition and adsorption energy. A stronger adsorption energy (more negative) suggests that the process is more thermodynamically favorable and is associated with a reduced power requirement to SET the device. Assuming a similar order of magnitude for the switching time across substrates (approximately tens to hundreds of nanoseconds[17, 18]), the switching power serves as an approximate measure of the switching energy.
Figure S2: Linear relationship between the computed adsorption energy of gold on various TMDs and the measured SET switching power, [14] indicating that stronger adsorption energies correlate with reduced switching power to SET the device. A linear fit is provided as a guide to eye on trend among the semiconducting materials; h-BN is omitted in the linear regression, as it is an wide band-gap insulator.
Figure S2 shows there is a correlation between the estimated measured switching power and the computed adsorption energy for various monolayer TMDs, lending credence to our model of adsorption energy as a proxy to switching power, and relatedly, switching energy. While adsorption energy is typically considered to be directly related to the RESET switching power in memristors—where the desorption and reformation of atomic or molecular states occur according to the DDA model—it is not solely related to the RESET process. The underlying reasons are more nuanced. One important factor is the asymmetry observed in the SET/RESET I-V characteristics of measured samples. This asymmetry is not fully captured by adsorption energies alone, suggesting that additional effects are at play during both switching processes. For instance, the Au electrode in MoS2/Au heterostructures induces strain on the MoS2 layer, leading to charge transfer and modifications in the electronic structure [34], which can affect the energetics of both the SET and RESET processes.
S2 Details on adsorption energetics
Figure S3: Adsorption energies calculated with the inclusion of van der Waals forces using a nonlocal vdW-DF functional as implemented in VASP.
Figure S3 shows the variation in adsorption energies with the consideration of van der Waals (vdW) forces and without. In the majority of cases, vdW corrections yield adsorption energies that are nearly identical to those calculated without considering vdW interactions. For group 10 and 11 metals (Ni, Pd, Pt, Cu, Ag, Au), there is a consistent shift of approximately 2.36 eV when vdW corrections are included. Nevertheless, the trends in adsorption energy are retained, which is the focus of our current study.
Figure S4: Adsorption energies of the different TM adsorbates on MoSe2, WS2, WSe2 (blue) and the change in Bader charge on the adatom (red) from the neutral state where negative values correspond with charge gain and positive values correspond with charge loss.
Figure S5: Adsorption energies of the different TM adsorbates on MoSe2, WS2, WSe2 (blue) using the chemical potential derived from a bulk adsorbate and the Bader charge (red)
Figure S6: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on MoSe2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S7: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on WS2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S8: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on WSe2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S9: Heat map of d-orbital contributions of the different adatoms onto a sulfur-vacant MoS2 and selenium-vacant MoSe2
Figure S10: Heat map of d-orbital contributions of the different adatoms onto a sulfur-vacant WS2 and selenium-vacant WSe2
Table S1: Adsorbates Exhibiting or Not Exhibiting Metallic-like Behavior on TMD Surfaces
Observed Behavior Adsorbates
Exhibits Enhanced Metallic-like Behavior Ag, Au, Cd, Co, Cr, Cu, Fe
Hf, Hg, Mn, Mo, Ni, Os, Pd
Pt Ru, Sc, Ta, Ti, W, Y, Zn
Does Not Exhibit Enhanced Metallic-like Behavior Rh, V, Zr, Nb, Re, Tc, Ir
S3 Additional SISSO details
In the construction of $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ we utilized basic algebraic operators given in Equation S1
$\hat{\mathbf{H}}^{(m)}\equiv\left\{I,+,-,\times,/,^{2},\sqrt{\text{ }},|-|\right\}$ (S1)
In addition, to prevent potential overfitting of data, we limit the maximal complexity of operators in a feature to 3.
Feature Name Abbreviation
1st Ionization Energy of adatom $IE_{\text{ads}}$
Electronegativity of adatom $X_{\text{ads}}$
Covalent Radius of Adatom $R_{c}^{\text{ads}}$
Weighted Workfunction of Adatom $\text{WF}_{\text{ads}}$
Electronegativity of cation $X_{\text{cat}}$
Covalent Radius of cation $R_{c}^{\text{cat}}$
Weighted Workfunction of cation $\text{WF}_{\text{cat}}$
Electronegativity of anion $X_{\text{an}}$
Covalent Radius of anion $R_{c}^{\text{an}}$
Weighted Workfunction of anion $\text{WF}_{\text{an}}$
Table S2: Primary features used in the SISSO-generated descriptor for adsorption energy. The table includes the electronegativity, covalent radius, and 1st ionization energy of adatoms, cations, and anions, as well as their weighted work functions. The electronegativity, covalent radius, and ionization energy values were obtained from [72], while the weighted work function values were obtained from [76]

### v2

@@ 2.4 Physical model for resistive switching based on point defects
As discussed in Section 1 of the Supplementary Information, a correlation between the measured switching energy [14] and the computed adsorption energies for the common TMDs suggests that the adsorption energy plays an important role in the energetics of resistive switching based on transition metal dichalcogenides. Previous studies, such as the one by Frey et al [43], have used a similar idea of using binding energies and formation energies to identify optimal defects for applications in neuromorphic and quantum information processing. Frey et al[43] used random forest regression to predict adsorption energies and identify potential descriptors but did not fully delve into the defect chemistry that dictated trends in adsorption energy. This study emphasizes understanding the adsorption mechanism between 2D TMDs and various transition metal adsorbates, to further elucidate how the valency, periodicity, and electronic configuration of the metal affect the adsorption energetics.
3 Discussion
@@
We have shown that the adsorption energy is relatively insensitive to the chemical species of the substrate when considering commonly studied TMDs (MoS2, MoSe2, WS2, WSe2). Saliently, the trends in adsorption species may be explained using the d-band model, a model based on electronic structure descriptors most well-known in its application to catalysts. Hence, the switching energy is materials-dependent and that dependence can be rationalized. An earlier study on the adsorption of Li in transition metal dichalcogenides for battery applications demonstrated that the work function can serve as a simple descriptor for explaining trends in Li adsorption[61]. In this study, we generalize this concept to adsorption energies of metal adatoms in low-dimensional materials. Thus, insights from electrochemistry and catalysis may be effectively cross-applied in resistive switching contexts, and can be applied more broadly to classes of substrates and adsorbates than previously demonstrated.
@@
Within and beyond resistive switching, defects in low-dimensional materials play a prominent role in next generation computing devices. Indeed, two-dimensional (2D) materials are ideal platforms for hosting quantum defects [63, 64]. An important implication of our study in rationalizing adsorption energies of point defect complexes in TMDs is the identification of electronic structure and atomic descriptors for both the substrate and electrode, which extends our ability to engineering designer defects for targeted applications. For practical device applications, we expect that the adsorption energy of transition metal adatoms plays a role in the switching energy (see also Supplementary Figure 2). That is, a large (negative) adsorption energy corresponds to a greater stability of the adsorbate, which may be important for long-term operation. On the other hand, a low (near zero) adsorption energy corresponds to a low switching energy, which may contribute to more energy-efficient resistive switching.
@@
Ultimately, we investigated the adsorption energies of various transition metal adsorbates on different monolayer TMD substrates, utilizing both computational methods and theoretical models to understand the underlying trends. We evaluated the adsorption energy of a metal adatom on a pre-existing anion vacancy for MoS2, MoSe2, WS2, and WSe2 and suggest these adsorption energies as proxy for investigating resistive switching energy. The adsorption energies displayed significant variation depending on the periodic group of the transition metal, with early transition metals (Hf, Zr, Ti) exhibiting the strongest adsorption energies, and late transition metals (Cu, Au, Ag, Zn, Cd, Hg) showing the weakest, while the middle transition metals exhibited moderate adsorption energies in between the early and late transition metals. We demonstrated that these trends in the adsorption energy could be largely captured by trends in the d-band center with respect to the Fermi level, and as a descriptor can capture most of the trend in adsorption energies. COHP analysis provided further insights into the nature of bonding and anti-bonding interactions between the adsorbates and their nearest neighbor sulfur atoms in the TMDs, emphasizing the importance of the local environment surrounding the adsorption site in determining the adsorption energy. We used this understanding to perform compressed sensing on atom-based descriptors, which does not require a DFT calculation. This SISSO analysis demonstrated promise in predicting absolute adsorption energies, using compound descriptors based off of the ionization energy, electronegativity, and atomic radii of the adsorbate and adsorption site.
Our findings highlight the importance of considering both the electronic structure and the local bonding environment in understanding materials-based trends in the adsorption behavior on TMD substrates to a greater degree than previously shown for transition metal adsorbate species. While the prediction of absolute adsorption defect energetics is on-going, we emphasize the understanding trends already provides materials design criteria for low-dimensional materials in energy-efficient resistive switching applications. Notably, this study provides a simplified materials-driven framework for modeling and understanding adsorption energy and resistive switching, achieved through cross-applying principles of the d-band model from electrocatalysis on adsorption mechanisms and energetics. By integrating multiple theoretical frameworks, including the d-band center model, bond analysis using COHP, and SISSO, we have presented a valuable tool for materials-driven design for future memristive devices.
4 Methods
4.1 Computational Parameters
The calculations in this work were performed using the Vienna Ab initio Simulation Package (VASP 6.3.0) [66, 67, 68] with the projector augmented-wave (PAW) method and the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [69, 70]. A kinetic energy cut-off of 400 eV was used for the plane-wave basis set expansion. The Brillouin zone was sampled using a 2 $\times$ 2 mesh, and spin polarization was taken into account, as it is expected to be significant when dealing with transition metal adsorbates [71]. Convergence criteria were set to an energy difference of $10^{-5}$ eV for electronic steps and force difference of 10 meV/Å. For density of states calculations, a $k$-point grid of 4 $\times$ 4 (equivalent to a resolution of 0.07 Å-1) was used, resulting in a difference less than 0.01 eV in Fermi energies).
To investigate the overall relationship between the structure and composition of a TMD-adatom complex, we use the adsorption energy defined by Eq. 1. Here, $E_{ads}$ is the adsorption energy, which represents the energy required for the adsorption of a single adatom onto the substrate. The terms $E^{tot}_{ads/sub}$, $E^{tot}_{sub}$, and $E^{tot}_{ads}$ denote the total energies of the structure with the adsorbate, the substrate without the adsorbate, and the chemical potential of the adatom, respectively.
$E_{ads}=E^{tot}_{ads/sub}-E^{tot}_{sub}-E^{tot}_{ads}$ (1)
In general, the adsorption energy depends on the chemical potential of the adatom, and hence depends on the bath from which the adatom originates, which in turn is determined by experimental conditions. We explore the effect of the chemical potential, ranging from the isolated adatom to that referenced to the bulk metal, on the adsorption energy in Section 2.1. In Eq. 1, we assume the chemical potential to be effectively that of an isolated adatom in order to focus on the adsorption energetics due to chemical and electronic differences in the adsorbate or monolayer. In general, other factors such as diffusion barriers are expected to influence memristive switching. Indeed, prior experimental and theoretical studies, including recent work from Anvari and Wang [34], have demonstrated that interfacial effects such as strain, interlayer spacing, and associated electronic modifications strongly impact the switching characteristics. Nevertheless, adsorption energies reflect energetically favorable initial and final configurations that are consistent with experimentally observed switching states [46]. Thus, our present analysis of adsorption energetics provides a foundation for ongoing and future studies exploring these dynamic processes in greater detail.
To minimize spurious defect-defect interactions, a supercell size of 8 × 8 (192 atoms) was used with a vacuum size of 20 Å. The relaxed structures shown in Figure 8 illustrate the distance between adjacent molybdenum atoms near the defect site. We focus on the adsorption of a metal adatom on a pre-existing chalcogen vacancy as the representative point defect complex, as our prior work has shown this configuration to be more energetically favorable than the adsorption of the adatom on a pristine monolayer [34]. For example, adsorption energies of Ag and Au on pristine MoS2 were calculated to be -0.59 eV and -0.64 eV, respectively, at the site directly above the molybdenum, with similar energies within 0.2 eV at other sites (-0.74 eV above the chalcogen site, -0.73 at the hollow site). In contrast, the adsorption energy at the chalcogen vacancy site was calculated to be -2.56 eV for gold. This is consistent with the dissociation-diffusion-adsorption (DDA) model [14] presented in Section 2.4, where the mechanism involves the dissociation of the adatom from the electrode, diffusion across the TMD, and eventual adsorption at an anion vacancy site.
4.2 The d-band model
Here, we apply the d-band model to rationalize trends in adsorption energy as they relate to resistive switching applications. The d-band model is based on the Newns-Anderson model [81, 82], describes the interaction between an adsorbate’s energy levels and the electronic structure of the substrate. This model provides insight into how the position and width of the substrate’s d-band influence adsorption strength and chemical reactivity. Subsequent refinements to this model have enhanced its accuracy and applicability, such as the inclusion of spin [71] and the inclusion of structural information via compound descriptors [62], and we utilize aspects of these refinements in our study. This approach utilizes the energy level of the d-band center and the shift induced by the adsorbate as key descriptors for adsorption energy.
[…]
where $\varepsilon$ represents the energy eigenvalues, and $n(\varepsilon)$ is the density of states of the bands with projected d-character. In our calculations, we define the d-band center over a fixed energy range that includes both regions below and above the Fermi level between the different adsorbates. This approach allows for a direct comparison of d-band centers across different adsorbates.
In this study, we select an energy range of $-10\,\mathrm{eV}$ to $1.5\,\mathrm{eV}$, referencing $E_{F}=0\,\mathrm{eV}$, to ensure that the d-band contributions both below and above the Fermi level are adequately captured. This range is chosen to ensure consistent comparison across different substrate and adsorbate configurations. To verify our choice of energy range, we computed the d‑band center using slightly extended and reduced windows (e.g., -11 eV to 1.5 eV and -10 eV to 2 eV) and observed that the calculated d‑band center values remained stable, confirming that our selected range reliably captures the relevant d‑state contributions.
Here, $f_{\sigma}$ represents the fractional occupancy of spin state $\sigma$, and $\varepsilon_{d\sigma}$ is the conventional Hammer-Nørskov d-band center for spin state $\sigma$. The final term, $\mu$, is the reduced fractional moment, defined as $\mu=\frac{f_{\uparrow}-f_{\downarrow}}{f_{\uparrow}+f_{\downarrow}}$. The first term in the equation gives the spin-averaged d-band center, where the spin-up and spin-down channels are averaged; the second term introduces a shift in the d-band center that depends on the spin polarization and the individual spin-resolved d-band centers. The orbital occupation and shift of the d-band center will be the two main parameters of focus in this study. Other parameters in the d-band model include the orbital occupation, the coupling between the electronic states of the adsorbate and substrate, and the energy associated with the orthogonalization of the states between the adsorbate and substrate. These descriptors are largely based on the electronic structure and thus suggest a generality of the d-band model beyond catalysis, i.e., for transition metal adsorbates, extending to applications such as resistive switching based on the formation of point defect complexes.
As a complement to the d-band center model, we also utilize crystal orbital Hamilton population analysis (COHP) to further study the bonding nature of the adsorbate on the various TMDs [84]. These calculations are performed using the Local-Orbital Basis Suite Towards Electronic-Structure Reconstruction (LOBSTER) package [85, 86, 87, 88, 89]. Crystal Orbital Hamilton Population (COHP) analysis distinguishes between bonding, non-bonding, and antibonding interactions between adjacent atoms by analyzing the projected density of states (DOS) weighted by Hamiltonian matrix elements. Specifically, the COHP between atomic orbitals $\phi_{\mu}$ and $\phi_{\nu}$ at energy $E$ is calculated as follows:
$\mathrm{COHP}_{\mu\nu}(E)=H_{\mu\nu}\sum_{j}c_{\mu j}c_{\nu j}\,\delta(E-\epsilon_{j})$
Here, $H_{\mu\nu}$ is the Hamiltonian matrix element between orbitals $\mu$ and $\nu$, $c_{\mu j}$ and $c_{\nu j}$ are the expansion coefficients of atomic orbitals in eigenstate $j$, and $\epsilon_{j}$ is the eigenvalue corresponding to eigenstate $j$. The sign of the Hamiltonian-weighted projected DOS (positive or negative) indicates the nature of the interactions: negative COHP values represent bonding (energy-lowering, stabilizing) interactions, whereas positive COHP values represent antibonding (energy-raising, destabilizing) interactions.
To quantify the bonding and antibonding interactions, we integrate the COHP over the energy range from $\infty$- ∞ to the Fermi level ($E_{F}$). The antibonding contribution corresponds to the integral over the positive part of COHP ($\mathrm{COHP}^{+}$):
[…]
Similarly, the bonding contribution corresponds to the integral over the negative part of COHP ($\mathrm{COHP}^{-}$):
[…]
The percent antibonding character is then calculated as:
[…]
This method provides a detailed understanding of chemical bonding by explicitly linking the DOS with the underlying electronic interactions dictated by the Hamiltonian, enabling a quantitative distinction between bonding and antibonding states across the relevant electronic structure of the system.
Finally, we use the Sure Independence Screening and Sparsifying Operator (SISSO) [54] to develop interpretable descriptors that leverage chemical intuition to predict adsorption energy (see Supplementary Table 3 in the SI for the complete list of descriptors). SISSO is a symbolic regression method that systematically combines chemically relevant primary features—such as ionization energy, electronegativity, and atomic radii—using mathematical operations to generate a vast space of potential descriptors. By creating a large pool of candidate features and applying a sparsifying operator, SISSO identifies a minimal set of the most predictive and interpretable descriptors, effectively capturing nonlinear relationships between input features and adsorption energy. This approach enables models that are physically meaningful and avoids overfitting by selecting descriptors that generalize well across different datasets. The strength of SISSO lies in its ability to reduce the problem’s dimensionality, distilling complex material behaviors into a few key descriptors. This enhances predictive accuracy and provides significant insights into the determinants of adsorption energy. Further details are provided in Supplementary Information (SI).
Data Availability
The main data supporting the findings of this study are provided within the paper and Supplementary material. Additional data is openly available on Zenodo at DOI: 10.5281/zenodo.13892108, Ref [90]
Acknowledgments
@@
Competing Interests
Authors B.H.L, J.F, W.W declare no financial or non-financial competing interests. Author D.A serves as an Editorial Board Member of this journal and had no role in the peer-review or decision to publish this manuscript. Author D.A declares no financial competing interests.